# A Comprehensive Survey of Artificial Intelligence Techniques for Talent Analytics

### v2

@@ 6.7 The Impact of Emergent Events on Talent Analytics
The emergent events like COVID-19 pandemic have had a huge impact on businesses and organizations, and there are many management scenarios where we can apply AI methods to effectively deal with these dilemmas. In talent management, first of all, because of the pandemic, the daily tasks of the employees in many companies will change, in order to make the company adapt to this fluctuation and to mitigate the impact of COVID-19, we can predict the performance of the employees by using AI methods [167]. Additionally, the epidemic has taken a toll on numerous markets and industries, causing employees in companies to face high psychological stress, and some may even leave the organization. We can complete the simulation of employees’ stress through the technology of artificial intelligence [134] and understand the factors that influence the employees’ intention to leave the organization [297, 389]. From the perspective of an organization’s leader, it is essential to understand the impact of the epidemic on customers so that an effective response can be taken. Therefore, completing sentiment analysis of customers through text is a very effective approach [70, 359]. Lastly, in the labor market, AI methods can effectively use large-scale data to provide researchers with research perspectives and technical support to study the impact of COVID-19 on the labor market [67, 11].
@@
Besides the organizational resilience, the forms of talent analytics are facing a transformation due to COVID-19. Because issues such as the need to limit social distances that arose during the epidemic triggered us to think about increasing human-computer interaction in the workplace scenario [291]. In the recruitment process, since global cooperation and telecommuting have become frequent, we can use artificial intelligence technologies such as natural language processing technology to screen job applicants in global real-time big data [379], in order to reduce the risk of disease transmission in densely populated scenarios, the use of chatbots to communicate with onboarding employees during the onboarding process can avoid the contact of the employees and simplify the process of onboarding employees [413]. Once employees are onboarded, they can also be trained with the help of a robot [397]. Xu et al. found that applying AI-enhanced VR simulators to employee training can reduce human resource costs and increase employee satisfaction, thereby enhancing the competitiveness of the organization [424]. Finally, for managers of organizations, as the temporary workforce of organizations is increasing after the epidemic era, combining temporary workforce management with AI can reduce risks and costs and improve work efficiency and quality [281].
7 Conclusions
Artificial Intelligence (AI)-driven talent analytics represents a potent frontier of innovation and opportunity in today’s competitive and fast-evolving business environment, even with the strong automation capability of LLM for decision making. However, in the process of the development of AI-based methods in talent analytics, many challenges need to be considered, such as fairness, explainability, data imbalance, temporal variation, etc. The solution of these problems can be more effective in evaluating AI, making AI-based decisions more accurate and usable. The essence of this survey provides a comprehensive exploration of the recent advancements in this domain. Specifically, we first delineated a detailed taxonomy of pertinent data, establishing a critical foundation for the utilization of AI techniques to understand talents, organizations, and management better. Subsequently, we illustrated the research efforts of the AI techniques for talent analytics, parsed into three critical aspects: talent management, organization management, and labor market analysis. Finally, we summarized the open challenges and potential prospects for future research directions within the AI-driven talent analytics sphere. The primary intent of this survey paper is to provide a thorough understanding of the recent effort of this emergent field to our readers, thereby fostering insights into the dynamic intersection of AI and talent analytics.

### v3

@@ C.7 The Impact of Emergent Events on Talent Analytics
Emergent events such as pandemics have had a huge impact on businesses and organizations, and there are many management scenarios where we can apply AI methods to effectively deal with these dilemmas. In talent management, because of the pandemic, the daily tasks of the employees in many companies will change. In order to make the company adapt to this fluctuation and to mitigate the impact of these emergent events, we can predict the performance of the employees by using AI methods [300]. Additionally, the epidemic has taken a toll on numerous markets and industries, causing employees in companies to face high psychological stress, and some may even leave the organization. We can complete the simulation of employees’ stress through AI technologies [402] and understand the factors that influence the employees’ intention to leave the organization [403, 404]. From the perspective of an organization’s leader, it is essential to understand the impact of the epidemic on customers so that an effective response can be taken. Therefore, completing sentiment analysis of customers using text is a very effective approach [301, 302]. Lastly, in the labor market, AI methods can effectively use large-scale data to provide researchers with research perspectives and technical support to study the impact of emergent events on the labor market [405, 406].
@@
Besides organizational resilience, the forms of talent analytics are facing a transformation due to some emergent events. Because issues such as the need to limit social distances that arose during the epidemic triggered us to think about increasing human-computer interaction in the workplace scenario [407]. In the recruitment process, since global cooperation and telecommuting have become frequent, we can use AI technologies such as natural language processing technology to screen job applicants in global real-time big data [303], to reduce the risk of disease transmission in densely populated scenarios, the use of chatbots to communicate with onboarding employees during the onboarding process can avoid the contact of the employees and simplify the process of onboarding employees [304]. Once employees are onboarded, they can also be trained with the help of a robot [305]. Xu et al. found that applying AI-enhanced VR simulators to employee training can reduce human resource costs and increase employee satisfaction, thereby enhancing the competitiveness of the organization [408]. Finally, for managers of organizations, as the temporary workforce of organizations is increasing after the epidemic era, combining temporary workforce management with AI can reduce risks and costs and improve work efficiency and quality [409].